a satellite orbits earth at an altitude 356 kilometers above the planets surface. Compute for its (a) radius (b) speed (c) orbital period. (mE=5.97x10^24 kg)


1
Expert's answer
2022-02-02T14:42:07-0500

Radius

"R=R_e+h"


"R=6400000+356000=6756000m"

We know that

In circular motion

"\frac{GmM}{R}=\frac{mv^2}{R}"

"V=\sqrt{\frac{GM}{R}}"

Put (R)=Re+h,M=Me

Speed "V=\sqrt{\frac{Gm_e}{R_e+h}}"

"V=\sqrt{\frac{6.67\times10^{-11}\times5.97\times10^{24}}{6400\times10^3+356000}}"

"V=\sqrt{\frac{6.67\times10^{-11}\times5.97\times10^{24}}{6756000}}=7.6772km/sec"

V=7.677km/sec

Time period

"T=\frac{2\pi R}{V}"

"T=\frac{2\times3.14\times6756000}{7677}=5526.59sec"

"T=\frac{5526.59}{60\times60}=1.53hr"